# The Racah algebra of rank 2: Properties, symmetries and representation

### v1

@@ 2.2 The rank-2 and higher Racah algebras
Proof: From the definition of the $P_{ij}$ and $P_{i}$ (2.16) and using the algebra decomposition relations (2.12) it is clear that the $P$’s generate the algebra and are symmetric under interchange of indices. Lemma 2 shows that the $D_{ijk}$ are antisymmetric under interchange of indices.
Commutation relation (2.31) is the definition of the $D$’s while relation (2.32) follows directly from the change of basis applied to (3.12). Lemma 6 gives relation (2.33) and Lemma 8 gives (2.34). Finally, the identity (2.35) follows from the discussion above under a cyclic permutation, or equivalently using the Jacobi identity with $[D_{jk\ell},[P_{ik},P_{k\ell}]]$.
It remains to consider Jacobi identities more generally. First, we begin with the case of three $P^{\prime}s$. The Jacobi identity for triples $\{P_{jk},P_{jk},P_{ij}\}$ and $\{P_{ki},P_{ij},P_{jk}\}$ are verified using Lemma 4. The case $\{P_{jk},P_{jk},P_{i\ell}\}$ is trivial since everything commutes. The final case is a triple of the form $\{P_{i\ell},P_{jk},P_{ij}\}$ as in
$[P_{i\ell},[P_{jk},P_{ij}]]+[P_{jk},[P_{ij},P_{i\ell}]]+[P_{ij},[P_{i\ell},P_{jk}]]$
$=2[P_{i\ell},D_{kji}]+2[P_{jk},D_{ji\ell}].$
Here we have use the definition of the $D$’s and the fact that $P$’s with disjoint indices commute. Lemma 7 gives
$[P_{\ell i},D_{ikj}]=[P_{kj},D_{j\ell i}]$
and so the remaining terms cancel.
For triples containing two $P$’s and a $D$, a first choice is for all three to belong to an $R(3)$ sub-algebra. This case was addressed in ther previous section, Lemma 1.
For the other two cases we assume that the $P$’s are not identical and have all 4 indices contained in them. Without loss of generality, we take $D_{ijk}$ and then either $P_{i\ell}$ and $P_{k\ell}$ or $P_{i\ell}$ and $P_{jk}.$ The first case was discussed above where the identity (2.35) was derived.
@@
$[P_{i\ell},[P_{jk},D_{ijk}]]=[P_{i\ell},\left(P_{jk}-2P_{j}\right)P_{ik}-P_{ij}\left(P_{jk}-2P_{k}\right)]$
$=2\left(P_{jk}-2P_{j}\right)D_{\ell ik}-2D_{\ell ij}\left(P_{jk}-2P_{k}\right).$
@@
Thus, we have shown that the Jacobi relations for two $P$’s and a $D$ are satisfied without additional constraints. The proof of higher-order relations follow similarly and are included in Appendix A.
Let us finish the section by collecting the results above that hold in $R(n)$ and compute the last commutation relation from the definition in Ref. [5], thus showing that our two definitions agree. We conjecture that in this case as well the Jacobi identity is satisfied without additional constraints, though we leave this for future investigation.
Theorem 2.
The Racah algebra $R(n)$ is generated by the elements $P_{i}$, $P_{ij}$ (2.16) and $D_{ijk}$ where the $P$’s are symmetric under interchange of the indices and the $D$’s are anti-symmetric. The generators satisfy the following commutation relations (for distinct indices)
$[P_{ij},P_{jk}]=2D_{ijk}$ (2.36)
@@
Proof: Relations (2.36-2.39) and (2.42) were shown for arbitrary indices. It remains to check the final relations (2.40) and (2.41) which have no analog when $n=4.$
For (2.40) we begin with the identity
@@ 3 Generators, presentations, and symmetries
As is apparent from the decomposition relations (2.3) in the rank-1 case and (2.12) for rank-2, the generator discussed above are not linearly independent. In this section, we will discuss possible choices of linear independent generators, in preparation for the following discussion of representations of the algebras. Different choices of basis will break permutation symmetry of the full algebra and we will briefly discuss what is left of the symmetry of the chosen basis.
@@ 3.1 The rank-1 Racah algebra
In the rank-1 case there are 4 central elements $C_{1}$, $C_{2},$ $C_{3}$ and $C_{123}.$ If we begin with those elements, then any choice of two of the remaining generators will be linearly independent and the third is obtained using (2.3). Any choice is equivalent up to permutation symmetry. A common choice is $C_{12},C_{23}$ as described in Figure 2. The symmetry of the generators can be realized by the interchange $1\longleftrightarrow 3$.
@@
Figure 1: $\mathcal{D}_{2}$ (or $\mathcal{P}_{2}$) symmetry illustration. The dashed lines link two generators that commute, which solid lines link two generators that do not commute.
@@ 3.2 The rank-2 Racah algebra
The rank-2 Racah algebra $R(4)$ is generated by 15 elements, that is 4 elements with one index ($C_{i}$), 6 elements with two indices ($C_{ij}$), 4 elements with three indices ($C_{ijk}$), and one maximal element $C_{1234}$. However, not all of these are linearly independent because of equation (2.12). In fact, it is possible to reduce the set of generators to 10 elements. It is convenient to keep the 5 central terms $C_{i}$ and $C_{1234}$, but various choices can be made for the non-central terms. Many authors are keeping 5 two-indices elements, which possesses a dihedral-group symmetry ($D_{4}$, isomorphic to the symmetries of a square). However, let us consider the contiguous basis, that is all the elements with continuous indices,
$C_{1},C_{2},C_{3},C_{4},\qquad C_{12},C_{23},C_{34},\qquad C_{123},C_{234},\qquad C_{1234}.$ (3.1)
One nice property of this choice is that a dihedral group symmetry $\mathcal{D}_{5}$ appears naturally. One can consider the following diagram.
@@
In addition to the dihedral group $\mathcal{D}_{5}$, the full algebra is invariant under the permutation group $\mathcal{P}_{4}$ acting on indices. While the $\mathcal{P}_{4}$ action also preserves the pentagon shape, the generators are not the the contiguous basis anymore (unless one considers the identity transformation). From here, it is interesting to assign more abstract names to those elements and consider any element that can be obtained by a $\mathcal{D}_{5}$ or $\mathcal{P}_{4}$ transformation. Let us assign up to $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$ transformation the following:
$\begin{array}[]{lll}\omega_{0}=C_{1234},&&\Omega_{0}=C_{23},\\ \omega_{1}=C_{1},&&\Omega_{1}=C_{34},\\ \omega_{2}=C_{2},&&\Omega_{2}=C_{123},\\ \omega_{3}=C_{3},&&\Omega_{3}=C_{234},\\ \omega_{4}=C_{4},&&\Omega_{4}=C_{12}.\end{array}$ (3.2)
@@
One can also introduce new notation $\Gamma_{i}$ for the commutators, what we called $D_{ijk}$. They are defined as follow:
$\Gamma_{i}=\dfrac{1}{2}[A_{i+2},A_{i-2}],\qquad i=0,...,4\qquad\mbox{mod }5.$ (3.3)
The notation may suggest that there are 5 commutators $\Sigma_{i}$, but they are linearly dependent, i.e.
$\Gamma_{0}+\Gamma_{1}+\Gamma_{2}+\Gamma_{3}+\Gamma_{4}=0.$ (3.4)
If one consider the group action of $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$, the sign of $\Sigma_{i}$ must be flipped depending on the group transformation. A rotation by $52n\pi/$ 5 from the dihedral group $\mathcal{D}_{5}$ act as follow
$\Omega_{i}\mapsto\Omega_{i+n}\qquad\omega_{i}\mapsto\omega_{i+n},\qquad\Gamma_{i}\mapsto\Gamma_{i+n},\qquad\mbox{mod }5,$ (3.5)
while an inversion with respect to a vertex $i$ ($A_{i}$) and its opposite edge ($\omega_{1}$) acts as follow
$\Omega_{i+j}\longleftrightarrow\Omega_{i-j},\qquad\omega_{i+j}\longleftrightarrow\omega_{i-j},\qquad\Gamma_{i+j}\longleftrightarrow-\Gamma_{i-j},\qquad\mbox{mod }5.$ (3.6)
Note that the mod 5 applies to the index, e.g. $\omega_{6}=\omega_{1}$.
@@
  • The definition of the 4 commutators:
    $[\Omega_{i+2},\Omega_{i-2}]=2\Gamma_{i},$ (3.7)
    $\sum_{i=0}^{4}\Gamma_{i}=0.$ (3.8)
  • The 5 central terms:
    $[\omega_{i},\omega_{j}]=[\omega_{i},\Omega_{j}]=[\omega_{i},\Gamma_{j}]=0.$ (3.9)
  • If two $\Omega$s are not next to one another on the pentagon, they commute:
@@
  • If an $\Omega$ and a $\Gamma$ share the same index, they commute:
    $[\Omega_{i},\Gamma_{i}]=0.$ (3.11)
  • The “inner” equations, i.e. the equations that can be used for $R(3)$-subalgebra:
    $~{}[\Omega_{i},\Gamma_{i+2}]$ $=$ $\Omega_{i}\Omega_{i+2}-\{\Omega_{i},\Omega_{i-1}\}-\Omega_{i}^{2}+(\omega_{i+1}+\omega_{i+2}+\omega_{i+3})\Omega_{i}$ (3.12)
    $+(\omega_{i+2}-\omega_{i+3})\Omega_{i+2}+\omega_{i+1}(\omega_{i+3}-\omega_{i+2}).$
  • The “outer” equations:
    $[\Omega_{i},\Gamma_{i+1}]$ $=$ $\sum_{k=0}^{4}\frac{(-1)^{k}}{2}\{\Omega_{i+k},\Omega_{i+k-1}\}+\Omega_{i}\Omega_{i+3}$ (3.13)
    $-\omega_{i+1}(\Omega_{i}+\Omega_{i+1})-\omega_{i+2}(\Omega_{i+2}+\Omega_{i+3})$
    $-\omega_{i-1}\Omega_{i-1}+\omega_{i+1}\omega_{i+2}+\omega_{i+1}\omega_{i-1}+\omega_{i+2}\omega_{i-1}.$
@@
Furthermore, the rank-2 Racah algebra possesses many rank-1 Racah subalgebras. Five of those subalgebras appear naturally when looking at Figure 2. The rank-1 Racah algebra made of $C_{12},C_{23},C_{123}$ and central terms has been highlighted in Figure 2, but any $\mathcal{D}_{5}$ action will also give a rank-1 Racah subalgebra, e.g. $\{C_{23},C_{34},C_{324},C_{2},C_{3},C_{4}\}$ or $\{C_{123},C_{234},C_{23},C_{1},C_{4},C_{1234}\}$.
The last observation in this section is regarding Casimirs of the rank-2 Racah algebra. The Casimir element of the rank-1 case (see equation (2.10)) can be written as
$\mathfrak{C}_{2}$ $=$ $\Gamma_{2}^{2}-\frac{1}{2}\{\Omega_{4}^{2},\Omega_{0}\}-\frac{1}{2}\{\Omega_{0}^{2},\Omega_{4}\}+\Omega_{4}^{2}+\Omega_{0}^{2}+\{\Omega_{4},\Omega_{0}\}$ (3.14)
$+\dfrac{1}{2}(\omega_{1}+\omega_{2}+\omega_{3}+\Omega_{2})(\{\Omega_{4},\Omega_{0}\}-2\Omega_{4}-2\Omega_{0})\}$
$+(\omega_{2}-\omega_{3})(\Omega_{2}-\omega_{1})\Omega_{4}+(\omega_{2}-\omega_{1})(\Omega_{2}-\omega_{3})\Omega_{0}$
$+(\omega_{1}+\omega_{3})(\Omega_{2}+\omega_{2})+(\omega_{1}\omega_{3}-\Omega_{2}\omega_{2})(\Omega_{2}-\omega_{1}+\omega_{2}-\omega_{3}).$
@@
$\mathfrak{C}_{i}$ $=$ $\Gamma_{i}^{2}-\frac{1}{2}\{\Omega_{i+2}^{2},\Omega_{i-2}\}-\frac{1}{2}\{\Omega_{i-2}^{2},\Omega_{i+2}\}+\Omega_{i+2}^{2}+\Omega_{i-2}^{2}+\{\Omega_{i+2},\Omega_{i-2}\}$ (3.15)
$+\dfrac{1}{2}(\omega_{i-1}+\omega_{i}+\omega_{i+1}+\Omega_{i})(\{\Omega_{i+2},\Omega_{i-2}\}-2\Omega_{i+2}-2\Omega_{i-2})\}$
@@
$+(\omega_{i-1}+\omega_{i+2})(\Omega_{i}+\omega_{i})$
$+(\omega_{i-1}\omega_{i+1}-\Omega_{i}\omega_{i})(\Omega_{i}-\omega_{i-1}+\omega_{i}-\omega_{i+1}),$
which can be obtained using the $\mathcal{D}_{5}$ symmetries, are Casimirs. This matches the results in [4, 10, 12]. Each Casimir $\mathfrak{C}_{i}$, $i=0,...,4$ corresponds to the Casimir of a rank-1 Racah subalgebra. As Crampé et al. noted in [4], it is quite surprising that the Casimir of a $R(3)$ subalgebra is also a Casimir of the whole algebra $R(4)$.

### v2

@@ 2.2 The rank-2 and higher Racah algebras
Proof.
From the definition of the $P_{ij}$ and $P_{i}$ (2.12) and using the algebra decomposition relations (2.9), it is clear that the $P$’s generate the algebra and are symmetric under interchange of indices. Lemma 2.4 shows that the $D_{ijk}$ are antisymmetric under interchange of indices.
Commutation relation (2.19) is the definition of the $D$’s while relation (2.20) follows directly from the change of basis applied to (3.2). Lemma 2.8 gives relation (2.21) and Lemma 2.10 gives (2.22). Finally, the identity (2.23) follows from the discussion above under a cyclic permutation, or equivalently using the Jacobi identity with $[D_{jk\ell},[P_{ik},P_{k\ell}]]$.
It remains to consider Jacobi identities more generally. First, we begin with the case of three $P$’s. The Jacobi identity for triples $\{P_{jk},P_{jk},P_{ij}\}$ and $\{P_{ki},P_{ij},P_{jk}\}$ are verified using Lemma 2.6. The case $\{P_{jk},P_{jk},P_{i\ell}\}$ is trivial since everything commutes. The final case is a triple of the form $\{P_{i\ell},P_{jk},P_{ij}\}$ as in
$[P_{i\ell},[P_{jk},P_{ij}]]+[P_{jk},[P_{ij},P_{i\ell}]]+[P_{ij},[P_{i\ell},P_{jk}]]=2[P_{i\ell},D_{kji}]+2[P_{jk},D_{ji\ell}].$
Here we have use the definition of the $D$’s and the fact that $P$’s with disjoint indices commute. Lemma 2.9 gives $[P_{\ell i},D_{ikj}]=[P_{kj},D_{j\ell i}]$ and so the remaining terms cancel.
For triples containing two $P$’s and a $D$, a first choice is for all three to belong to an $R(3)$ sub-algebra. This case was addressed in the previous section, Lemma 2.2.
For the other two cases, we assume that the $P$’s are not identical and have all 4 indices contained in them. Without loss of generality, we take $D_{ijk}$ and then either $P_{i\ell}$ and $P_{k\ell}$ or $P_{i\ell}$ and $P_{jk}$. The first case was discussed above where the identity (2.23) was derived.
@@
$[P_{i\ell},[P_{jk},D_{ijk}]]$ $=[P_{i\ell},(P_{jk}-2P_{j})P_{ik}-P_{ij}(P_{jk}-2P_{k})]$
$=2(P_{jk}-2P_{j})D_{\ell ik}-2D_{\ell ij}(P_{jk}-2P_{k}).$
@@
Thus, we have shown that the Jacobi relations for two $P$’s and a $D$ are satisfied without additional constraints. The proof of higher-order relations follow similarly and are included in Appendix A. ∎
Let us finish the section by collecting the results above that hold in $R(n)$ and compute the last commutation relation from the definition in [4], thus showing that our two definitions agree. We conjecture that in this case as well the Jacobi identity is satisfied without additional constraints, though we leave this for future investigation.
Theorem 2.12.
The Racah algebra $R(n)$ is generated by the elements $P_{i}$, $P_{ij}$ (2.12) and $D_{ijk}$ where the $P$’s are symmetric under interchange of the indices and the $D$’s are anti-symmetric. The generators satisfy the following commutation relations ((((for distinct indices)))):
$[P_{ij},P_{jk}]=2D_{ijk},$ (2.24)
$[P_{jk},D_{ijk}]=(P_{jk}-2P_{j})P_{ki}-P_{ij}(P_{jk}-2P_{k}),$ (2.25)
@@
Proof.
Relations (2.24)–(2.27) and (2.30) were shown for arbitrary indices. It remains to check the final relations (2.28) and (2.29) which have no analog when $n=4$.
For (2.28), we begin with the identity
@@ 3 Generators, presentations, and symmetries
As is apparent from the decomposition relations (2.2) in the rank-1 case and (2.9) for rank-2, the generators discussed above are not linearly independent. In this section, we will discuss possible choices of linear independent generators, in preparation for the following discussion of representations of the algebras. Different choices of basis will break permutation symmetry of the full algebra and we will briefly discuss what is left of the symmetry of the chosen basis.
@@ 3.1 The rank-1 Racah algebra
In the rank-1 case, there are 4 central elements $C_{1}$, $C_{2}$, $C_{3}$ and $C_{123}$. If we begin with those elements, then any choice of two of the remaining generators will be linearly independent and the third is obtained using (2.2). Any choice is equivalent up to permutation symmetry. A common choice is $C_{12}$, $C_{23}$ as described in Figure 1. The symmetry of the generators can be realized by the interchange $1\longleftrightarrow 3$.
@@
Figure 1: $\mathcal{D}_{2}$ (or $\mathcal{P}_{2}$, the permutation group of two elements) symmetry illustration, i.e., the dihedral group of a line. The dashed lines link two generators that commute, which solid lines link two generators that do not commute.
@@ 3.2 The rank-2 Racah algebra
The rank-2 Racah algebra $R(4)$ is generated by 15 elements, that is 4 elements with one index $(C_{i})$, 6 elements with two indices $(C_{ij})$, 4 elements with three indices $(C_{ijk})$, and one maximal element $C_{1234}$. However, not all of these are linearly independent because of equation (2.9). In fact, it is possible to reduce the set of generators to 10 elements. It is convenient to keep the 5 central terms $C_{i}$ and $C_{1234}$, but various choices can be made for the non-central terms. Many authors are keeping 5 two-indices elements, which possesses a dihedral-group symmetry ($\mathcal{D}_{4}$, isomorphic to the symmetries of a square). However, let us consider the contiguous basis, that is all the elements with continuous indices,
$C_{1},C_{2},C_{3},C_{4},\qquad C_{12},C_{23},C_{34},\qquad C_{123},C_{234},\qquad C_{1234}.$
One nice property of this choice is that a dihedral group symmetry $\mathcal{D}_{5}$ appears naturally. One can consider the diagram (see Figure 2).
[…]
@@
In addition to the dihedral group $\mathcal{D}_{5}$, the full algebra is invariant under the permutation group $\mathcal{P}_{4}$ acting on indices. While the $\mathcal{P}_{4}$ action also preserves the pentagon shape, the generators are not the contiguous basis anymore (unless one considers the identity transformation). It is interesting to note that if one combines $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$, one obtains a larger symmetry group isomorphic to $\mathcal{P}_{5}$. This larger symmetry group appears in [3] and is presented differently, expressing the action of $\mathcal{P}_{5}$ on the icosidodecahedron. However, we are interested here in investigating the symmetries of the particular choice of generators and the expression of the algebra relations in terms of these generators only, so we do not carry the entire symmetry group. Hence, we focus on $\mathcal{D}_{5}$ and a natural expression of its action on the contiguous basis, the pentagon.
From here, it is interesting to assign more abstract names to those elements and consider any element that can be obtained by a $\mathcal{D}_{5}$ or $\mathcal{P}_{4}$ transformation. Let us assign up to $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$ transformation the following:
$\omega_{0}=C_{1234},\qquad$ $\omega_{1}=C_{1},\qquad$ $\omega_{2}=C_{2},\qquad$ $\omega_{3}=C_{3},\qquad$ $\omega_{4}=C_{4},$
$\Omega_{0}=C_{23},\qquad$ $\Omega_{1}=C_{34},\qquad$ $\Omega_{2}=C_{123},\qquad$ $\Omega_{3}=C_{234},\qquad$ $\Omega_{4}=C_{12}.$
@@
One can also introduce new notation $\Gamma_{i}$ for the commutators, what we called $D_{ijk}$. They are defined as follows:
$\Gamma_{i}=\tfrac{1}{2}[\Omega_{i+2},\Omega_{i-2}],\qquad i=0,\dots,4,\qquad\mbox{mod }5.$
The notation may suggest that there are 5 commutators $\Gamma_{i}$, but they are linearly dependent, i.e., $\Gamma_{0}+\Gamma_{1}+\Gamma_{2}+\Gamma_{3}+\Gamma_{4}=0$. If one consider the group action of $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$, the sign of $\Gamma_{i}$ must be flipped depending on the group transformation. A rotation by $52n\pi/$ 5 from the dihedral group $\mathcal{D}_{5}$ acts as follows:
$\Omega_{i}\mapsto\Omega_{i+n}\qquad\omega_{i}\mapsto\omega_{i+n},\qquad\Gamma_{i}\mapsto\Gamma_{i+n},\qquad\mbox{mod }5,$
while an inversion with respect to a vertex $i$ ($\Omega_{i}$) and its opposite edge ($\omega_{1}$) acts as follows:
$\Omega_{i+j}\longleftrightarrow\Omega_{i-j},\qquad\omega_{i+j}\longleftrightarrow\omega_{i-j},\qquad\Gamma_{i+j}\longleftrightarrow-\Gamma_{i-j},\qquad\mbox{mod }5.$
Note that the mod 5 applies to the index, e.g., $\omega_{6}=\omega_{1}$.
@@
  • The definition of the $4$ commutators
    $[\Omega_{i+2},\Omega_{i-2}]=2\Gamma_{i},\qquad\sum_{i=0}^{4}\Gamma_{i}=0.$
  • The $5$ central terms $[\omega_{i},\omega_{j}]=[\omega_{i},\Omega_{j}]=[\omega_{i},\Gamma_{j}]=0$.
  • If two $\Omega$s are not next to one another on the pentagon, they commute
@@
  • If an $\Omega$ and a $\Gamma$ share the same index, they commute $[\Omega_{i},\Gamma_{i}]=0$.
  • The “inner” equations, i.e., the equations that can be used for $R(3)$-subalgebra
    $[\Omega_{i},\Gamma_{i+2}]=\Omega_{i}\Omega_{i+2}-\{\Omega_{i},\Omega_{i-1}\}-\Omega_{i}^{2}+(\omega_{i+1}+\omega_{i+2}+\omega_{i+3})\Omega_{i}$
    ${}+(\omega_{i+2}-\omega_{i+3})\Omega_{i+2}+\omega_{i+1}(\omega_{i+3}-\omega_{i+2}).$ (3.2)
  • The “outer” equations
    $[\Omega_{i},\Gamma_{i+1}]=\sum_{k=0}^{4}\frac{(-1)^{k}}{2}\{\Omega_{i+k},\Omega_{i+k-1}\}+\Omega_{i}\Omega_{i+3}-\omega_{i+1}(\Omega_{i}+\Omega_{i+1})$
    ${}-\omega_{i+2}(\Omega_{i+2}+\Omega_{i+3})-\omega_{i-1}\Omega_{i-1}+\omega_{i+1}\omega_{i+2}$
    ${}+\omega_{i+1}\omega_{i-1}+\omega_{i+2}\omega_{i-1}.$ (3.3)
@@
[…]
The last observation in this section is regarding Casimir operators of the rank-2 Racah algebra. The Casimir element of the rank-1 case (see equation (2.7)) can be written as
$\mathfrak{C}_{2}=\Gamma_{2}^{2}-\tfrac{1}{2}\big{\{}\Omega_{4}^{2},\Omega_{0}\big{\}}-\tfrac{1}{2}\big{\{}\Omega_{0}^{2},\Omega_{4}\big{\}}+\Omega_{4}^{2}+\Omega_{0}^{2}+\{\Omega_{4},\Omega_{0}\}$
${}+\tfrac{1}{2}(\omega_{1}+\omega_{2}+\omega_{3}+\Omega_{2})(\{\Omega_{4},\Omega_{0}\}-2\Omega_{4}-2\Omega_{0})+(\omega_{2}-\omega_{3})(\Omega_{2}-\omega_{1})\Omega_{4}$
${}+(\omega_{2}-\omega_{1})(\Omega_{2}-\omega_{3})\Omega_{0}+(\omega_{1}+\omega_{3})(\Omega_{2}+\omega_{2})$
${}+(\omega_{1}\omega_{3}-\Omega_{2}\omega_{2})(\Omega_{2}-\omega_{1}+\omega_{2}-\omega_{3}).$
@@
$\mathfrak{C}_{i}=\Gamma_{i}^{2}-\tfrac{1}{2}\big{\{}\Omega_{i+2}^{2},\Omega_{i-2}\big{\}}-\tfrac{1}{2}\big{\{}\Omega_{i-2}^{2},\Omega_{i+2}\big{\}}+\Omega_{i+2}^{2}+\Omega_{i-2}^{2}+\{\Omega_{i+2},\Omega_{i-2}\}$
${}+\tfrac{1}{2}(\omega_{i-1}+\omega_{i}+\omega_{i+1}+\Omega_{i})(\{\Omega_{i+2},\Omega_{i-2}\}-2\Omega_{i+2}-2\Omega_{i-2})$
@@
${}+(\omega_{i-1}+\omega_{i+1})(\Omega_{i}+\omega_{i})+(\omega_{i-1}\omega_{i+1}-\Omega_{i}\omega_{i})(\Omega_{i}-\omega_{i-1}+\omega_{i}-\omega_{i+1}),$
which can be obtained using the $\mathcal{D}_{5}$ symmetries, are Casimir operators. This matches the results in [3, 7, 8]. Each Casimir operator $\mathfrak{C}_{i}$, $i=0,\dots,4$ corresponds to the Casimir operator of a rank-1 Racah subalgebra. As Crampé et al. noted in [3], it is quite surprising that the Casimir operator of a $R(3)$ subalgebra is also a Casimir operator of the whole algebra $R(4)$.